# Optimizing the interaction geometry of inverse Compton scattering x-ray sources

### v1

@@ IV Conclusion
We have developed an analytical framework that optimizes the interaction geometry of an ICS x-ray source. In a head-on geometry a tight symmetrical laser pulse focus is required. The optimized laser spot size is primarily determined by the divergence of a focused laser pulse and the nonlinear broadening. In a grazing angle geometry we have found that an elliptical focus of the laser pulse is required to mitigate the reduced interaction time. Additionally the effects of laser pulse divergence and a three-dimensional electron bunch have been analyzed, resulting in a closed form expressions that allows the direct calculation of the laser pulse dimensions, interaction angle, and electron energy that optimize the x-ray brilliance. We find that this grazing angle geometry is especially useful for soft x-ray generation, a regime that is considered to be inefficient with conventional head-on scattering.
@@ Acknowledgements.
We thank Brian Schaap for various fruitful discussions. This research was funded by Interreg Vlaanderen-Nederland (Smart*Light project 2.0).
@@ Appendix A Derivation of the proper interaction time
with $\phi(\tau)=k_{L}^{\nu}x_{\nu}(\tau)$ the laser phase and $\epsilon^{\mu}$ is the laser 4-polarization. It is important to note that $A^{\mu}$ describes the vector potential as experienced by the electron, i.e. evaluated at the position and time described by the electron trajectory $x^{\mu}$. To arrive at Eq. (22) we assume that the amplitude vector potential remains small $(A_{0}\ll 1)$ and that the laser pulse envelope varies slowly compared to the laser phase. Additionally, we assume that the laser polarization is perpendicular to the initial electron 4-velocity $u_{0}^{\mu}$. The spectral angular density of the radiated field is given by [schaap2022photon]
@@
with $\varphi_{\mathrm{x}}$ the polar angle of the emitted radiation. In the exponent we can directly identify the angular dependent Doppler shift of Eq. (1) and the Fourier limited bandwidth $\Delta\omega/\omega_{\mathrm{x}}=\left[\sqrt{2}\gamma(1-\beta\cos{\theta_{L}})\omega_{L}\tau_{\mathrm{int}}\right]^{-1}$. Evaluation of Eq. (48) at $\omega=\omega_{\mathrm{x}}$ and $\theta_{x}=0$ results in Eq. (22). By integration over the x-ray frequency we find the angular radiation distribution
@@
Using $u_{0}^{\nu}r_{\nu}=\gamma\beta\sin{\theta_{L}}$ and $u_{0}^{\nu}k_{L\nu}=\gamma k_{L}(1+\beta\cos{\theta_{L}})$ the expression for the proper interaction time (Eq. (24)) can be identified. To incorporate the finite electron beam waist, we take $x_{0}^{\mu}=({0,x_{0},y_{0},z_{0}})$ and integrate over the electron distribution.
@@ Appendix B Comparison with simulations
Throughout this paper, the developed theory has been checked with simulations. This is done by numerically calculating the spectral angular density from the electron motion with Liénard-Wiechert potentials using the algorithm developed in [thomas2010algorithm]. The electron motion is calculated using a particle tracking simulation (General Particle Tracer [PulsarPhysicsGeneralGPT]) and the electro-magnetic field descriptions of a pulsed Gaussian $\mathrm{TEM}_{00}$ mode with a focus that can be asymmetrical.
@@ Appendix C Definition of x-ray brilliance
Throughout this paper we use the convention of r.m.s. x-ray brilliance, where $\Delta A=\pi\langle x^{2}+y^{2}\rangle$, $\Delta\Omega=\pi\langle\theta_{\mathrm{x}}^{2}\rangle$, and $\Delta\omega/\omega_{\mathrm{x}}=\sqrt{\langle\omega^{2}\rangle-\langle\omega\rangle^{2}}/\omega_{\mathrm{x}}$ refer to r.m.s. quantities. Since we assume a Gaussian electron beam, $\Delta A=2\pi\sigma_{e}^{2}$. However, since the radiation distribution to be approximately a uniform disk, $\Delta\Omega=\pi\Theta_{\mathrm{x}}^{2}/2$. To convert the results from this paper to a FWHM convention, one can use $B_{\mathrm{x}}^{FWHM}=B_{\mathrm{x}}/(2\sqrt{2\ln{2}})^{3}$.

### v2

@@ IV Conclusion
In this paper, we have presented a theoretical approach to optimize inverse Compton scattering x-ray sources. We have found closed form analytical expressions to describe the x-ray beam properties in terms of the laser and electron beam parameters. We find that the interaction angle is an important optimization parameter, which has not yet been treated in great detail. The model not only provides physical intuition on the important scaling of parameters involved in Compton scattering x-ray generation, but also allows for analytical optimization, which is usually done by time consuming simulations. The presented method can serve as a tool for the design of future Compton scattering x-ray sources and the optimization of existing inverse Compton scattering beamlines.
We find that, apart from the conventional head-on interaction geometry, a co-propagating grazing angle interaction can be particularly effective for generating soft x-rays. In general, the optimization involves a subtle interplay between several beam parameters. For a head-on geometry we find the surprising result that lowering the bunch charge can lead to a higher x-ray brilliance while maintaining the same x-ray flux. For soft x-ray wavelengths, optimization of a co-propagating grazing angle geometry results in an order of magnitude higher x-ray brilliance, compared to the optimized head-on geometry.
Even higher x-ray brilliance could be achieved in a head-on geometry by using chirped laser pulses to suppress the nonlinear broadening. The grazing angle geometry could be optimized even further by employing a laser pulse front tilt, which increases the interaction time without loss of intensity.
@@ Acknowledgements.
We thank Brian Schaap for various fruitful discussions. This research was funded by EU Interreg Vlaanderen-Nederland (Smart*Light project 2.0), by the Ministry of Economic Affairs in the Netherlands through a TKI-grant, and by the Dutch Research Council (NWO) in The Netherlands, through the Industrial Partnership Program (IPP) grant ’ColdLight: From laser-cooled atoms to coherent soft X-rays’.
@@ Appendix A Derivation of the proper interaction time
with $\phi(\tau)=k_{L}^{\nu}x_{\nu}(\tau)$ the laser phase and $\epsilon^{\mu}$ is the laser 4-polarization. It is important to note that $A^{\mu}$ describes the vector potential as experienced by the electron, i.e. evaluated at the position and time described by the electron trajectory $x^{\mu}$. To arrive at Eq. (22) we assume that the amplitude vector potential remains small $(A_{0}\ll 1)$ and that the laser pulse envelope varies slowly compared to the laser phase. Additionally, we assume that the laser polarization is perpendicular to the initial electron 4-velocity $u_{0}^{\mu}$. The spectral angular density of the radiated field is given by [37]
@@
with $\varphi_{\mathrm{x}}$ the polar angle of the emitted radiation. In the exponent we can directly identify the angular dependent Doppler shift of Eq. (2) and the Fourier limited bandwidth $\sigma_{\omega}/\omega_{\mathrm{x}}=\left[\sqrt{2}\gamma(1-\beta\cos{\theta_{L}})\omega_{L}\tau_{\mathrm{int}}\right]^{-1}$. Evaluation of Eq. (48) at $\omega=\omega_{\mathrm{x}}$ and $\theta_{x}=0$ results in Eq. (22). By integration over the x-ray frequency we find the angular radiation distribution
@@
Using $u_{0}^{\nu}r_{\nu}=\gamma\beta\sin{\theta_{L}}$ and $u_{0}^{\nu}k_{L\nu}=\gamma k_{L}(1-\beta\cos{\theta_{L}})$ the expression for the proper interaction time (Eq. (24)) can be identified. To incorporate the finite electron beam waist, we take $x_{0}^{\mu}=({0,x_{0},y_{0},z_{0}})$ and integrate over the electron distribution.
@@ Appendix B Comparison with simulations
Throughout this paper, the developed theory has been checked with simulations. This is done by numerically calculating the spectral angular density from the electron motion with Liénard-Wiechert potentials using the algorithm developed in [41]. The electron motion is calculated using a particle tracking simulation (General Particle Tracer [32]) and the electro-magnetic field descriptions of a pulsed Gaussian $\mathrm{TEM}_{00}$ mode with a focus that can be asymmetrical.
@@ Appendix C Definition of x-ray brilliance
Throughout this paper we define the x-ray brilliance such that
$\int B_{\mathrm{x}}\ \partial A\ \partial\Omega\ \partial(\omega/\omega_{\mathrm{x}})=fN_{\mathrm{x}},$ (54)
according to the convention used in [5]. Note that, using this convention, $B_{\mathrm{x}}$ is independent on whether r.m.s., FWHM or other values are used to the define the quantities $\Delta A,\Delta\Omega$ and $\Delta\omega/\omega_{\mathrm{x}}$, as long as a proper normalization is used. Additionally, this definition of x-ray brilliance ensures that Eq. 20 holds. Since $\sigma_{e}$ and $\Delta\omega/\omega_{\mathrm{x}}$ refer to r.m.s. quantities, $\Delta A=2\pi\sigma_{e}^{2}$ and $\Delta\omega/\omega_{\mathrm{x}}=\sqrt{2\pi}\sigma_{\omega}/\omega_{\mathrm{x}}$. Because the x-ray angular spread is uniform, $\Delta\Omega=\pi\Theta_{\mathrm{x}}^{2}$. Combining these expressions results in Eq. (1).